# Proximal Interacting Particle Langevin Algorithms

### v2

@@ Results.
We have analysed the sparsity of the weights from the final particle cloud for MYIPLA. Figure 5 shows how the performance metrics evolve when weights below a certain threshold are set to zero, when particles are averaged before (5(a)) or after (5(b)) computing the performance.
@@
Once we have set the weights of the matrix below a certain threshold to zero, it is necessary to explore the dead units. These are hidden units all of whose input or output weights are zero [Williams, 1995]. In both cases, the unit is redundant and it can be eliminated to obtain a functionally equivalent network architecture, we will called this new effective weight matrix $w_{\text{pruned}}$.
The occupancy ratio of a weight matrix $w$ [Marinó et al., 2023] is defined as $\psi=\text{size}(w_{\text{pruned}})/\text{size}(w)$, where size denotes the memory size. The inverse of $\psi$ is the compression ratio. We compute the occupancy ratio of the weight matrix for both the hidden and output layer for different values of the pruning threshold. We do this for each particle of the final cloud and obtain the average as well as for the averaged final particle cloud, results are shown in Figure 6.
@@
D.4 Sparse matrix completion
D.4.1 Example 1
@@
To obtain each of the matrices $X^{l}$ we generate two integers $a,b\in[-3,3]$, then
$X_{i,j}^{l}=\begin{cases}a+b+\varepsilon&\text{if }i\bmod 2=0,j\bmod 2=0\\ -a+b+\varepsilon&\text{if }i\bmod 2=1,j\bmod 2=0\\ a-b+\varepsilon&\text{if }i\bmod 2=0,j\bmod 2=1\\ -a-b+\varepsilon&\text{if }i\bmod 2=1,j\bmod 2=1\end{cases}$
where $\varepsilon\sim N(0,10^{-2})$. Each entry from the resulting $Y^{l}$ is observed with a probability of $0.5$.
@@
Recall that $g_{2}(\theta,\Phi,X)$ is of the form
$g_{2}(\theta,\Phi,X)=-\log(C(\theta))+e^{-\theta}\sum_{l}\|X^{l}\|_{\text{tr}}.$
Due to the difficulty of calculating $C(\theta)$ we assume that it is approximately constant $C(\theta)\approx C$. Thus, it is not necessary to take it into account when deriving the proximal operator.
To compute the proximal map, we first observe that if $\theta,\Phi$ are known, then by Cai et al. [2010] (Theorem 2.1)
$\operatorname{prox}_{g_{2}}^{\lambda}(X^{l})=\operatorname*{arg\,min}_{Z}\;\{e^{-\theta}\|Z\|_{\text{tr}}+\frac{1}{2\lambda}\|X^{l}-Z\|_{F}^{2}\}=S_{e^{-\theta}\lambda}(X^{l}):=U\Sigma_{e^{-\theta}\lambda}V^{T},$
where $X=U\Sigma V^{T}$ is the SVD, and $\Sigma_{\beta}$ is diagonal with entries $(\Sigma_{\beta})_{ii}=\max\{\Sigma_{ii}-\beta,0\}$. Based on this, we calculate
$\operatorname{prox}_{g_{2}}^{\lambda}(\theta,\Phi,X)=\operatorname*{arg\,min}_{(\alpha,\Gamma,Z)}\;\{e^{-\alpha}\sum_{l}\|Z^{l}\|_{\text{tr}}+\frac{1}{2\lambda}\big{(}\|\theta-\alpha\|^{2}+\|\Phi-\Gamma\|_{F}^{2}+\sum_{k}\|X^{l}-Z^{l}\|_{F}^{2}\big{)}\}.$
The minimisers $(\alpha,\Gamma,Z)$ satisfy the following system of equations
$\Gamma=\Phi,$
$\alpha=\theta+\lambda e^{-\alpha}\sum_{l}\|S_{e^{-\alpha}\lambda}(X^{l})\|_{\text{tr}}\Longrightarrow(\alpha-\theta)e^{\alpha-\theta}=\lambda e^{-\theta}\sum_{l}\|S_{e^{-\alpha}\lambda}(X^{l})\|_{\text{tr}},$ (56)
$Z^{l}=S_{\lambda e^{-\alpha}}(X^{l}).$ (57)
@@
$(\alpha-\theta)e^{\alpha-\theta}\approx\lambda e^{-\theta}\sum_{l}\|S_{e^{-\theta}\lambda}(X^{l})\|_{\text{tr}}\Longrightarrow\alpha\approx\theta+W(\lambda e^{-\theta}\sum_{l}\|S_{e^{-\theta}\lambda}(X^{l})\|_{\text{tr}}).$
Substituting this value of $\alpha$ into (57), we obtain
$Z^{l}\approx S_{e^{-\alpha}\lambda}(X^{l}).$
@@
To stabilise the implementation of MYIPLA, we have divided the gradients and proximal mapping term of the updates of $\theta$ and $\Phi$ by the dimension of the tensor $(Y^{1},\dots,Y^{L})$, that is, $m\times n\times L$. We then set $\gamma=0.05$ and $\lambda=0.35$.

### v3

@@ Results.
First, it is important to discuss whether the Laplace prior is more appropriate in this setting than the Normal one. Jaynes [1968] provides two reasons why the Laplace prior is particularly suitable for Bayesian neural network models. Firstly, for any feedforward network there is a functionally equivalent network in which the weight of a non-direct connection has the same size but opposite sign, therefore consistency demands that the prior for a given weight $w$ is a function of $|w|$ alone. Secondly, if it is assumed that all that is known about $|w|$ is its scale, and that the scale of a positive quantity is determined by its mean rather than some higher order moment, then the maximum entropy distribution for a positive quantity constrained to a given mean is the exponential distribution. It would follow that the signed weight $w$ has a Laplace density [Williams, 1995].
We have examined the sparsity-inducing nature of the Laplace prior versus a normal one in Figure 2 and Table 3. As mentioned in the main text, the sparse representation of our experiment also has the advantage of producing models that are smaller in terms of memory usage when small weights are zeroed out. To investigate this, we set to zero all weights below a certain threshold and analyse the performance of the compressed weight matrices. We consider two cases, averaging the particles of the final cloud $X_{500}^{1},\dots,X_{500}^{100}$, applying the threshold and then calculating the performance, and secondly, setting to zero small values of each particle of the cloud and averaging the performance of each particle. We compare the results for the Bayesian neural networks with Laplace and Normal priors (Table 8). It is important to note that when applying the same threshold to both cases, the Laplace prior leads to a very compressed weight matrix compared to the Normal prior, i.e. there is a significant difference in the percentage of weights set to zero. We observe that when setting the same proportion of weights to zero in both layers, the performance of the BNN with Laplace priors is better in terms of the log pointwise predictive density than that of the BNN with Normal priors, especially when averaging the final cloud of particles before computing the performance.
[…]
particles? Layer 1 Layer 2 Layer 1 Layer 2
Laplace $74$ $48$ $0.2$ $0.2$ $\mathbf{7.0}$ $\mathbf{-0.23}$
$56$ $35$ $1$ $1$ $\mathbf{1.5}$ $\mathbf{-0.07}$
Normal $74$ $48$ $0.5$ $1.1$ $15$ $-0.74$
$16$ $15$ $0.2$ $0.2$ $16$ $-0.78$
$56$ $35$ $7$ $4$ $2.0$ $-0.11$
$8.6$ $7.1$ $1$ $1$ $1.5$ $-0.10$
Figure 6 shows how the performance metrics evolve when weights below a certain threshold are set to zero, when particles are averaged before (6(a)) or after (6(b)) computing the performance for MYIPLA.
Once we have set the weights of the matrix below a certain threshold to zero, it is necessary to explore the dead units. These are hidden units all of whose input or output weights are zero [Williams, 1995]. In both cases, the unit is redundant and it can be eliminated to obtain a functionally equivalent network architecture, we will called this new effective weight matrix $w_{\text{pruned}}$. The occupancy ratio of a weight matrix $w$ [Marinó et al., 2023] is defined as $\psi=\text{size}(w_{\text{pruned}})/\text{size}(w)$, where size denotes the memory size. The inverse of $\psi$ is the compression ratio. We compute the occupancy ratio of the weight matrix for both the hidden and output layer for different values of the pruning threshold. We do this for each particle of the final cloud and obtain the average as well as for the averaged final particle cloud, results are shown in Figure 7.
@@
E.3.2 Sparsity Inducing Prior: CIFAR10
We further evaluate our methods on a classification task using a more complex dataset: CIFAR10. As with the MNIST dataset, to reduce the cost of computing the gradients on a big dataset, we subsample 5000 data points with labels plane, car, ship and truck.
Given that the data consists of colour images, we employ a convolutional neural network (CNN) architecture. Specifically, we use a combination of convolutional layers, max pooling layers, and linear layers with non-linear activation functions. For simplicity, we apply a sparsity-inducing prior only to the linear layers, and not to the convolutional ones. The sparsity inducing prior for each layer with weight matrix $w$ is given by $p_{\alpha}(w)=\prod_{i}\text{Laplace}(w_{i}|0,e^{2\alpha})$ where $\alpha$ is learn from the data. The network structure and layer dimensions are as follows.
  • Convolutional layer (Deterministic): Conv2d(3, 6, 5)
  • Max pooling layer (Deterministic): MaxPool2d(2, 2)
  • Convolutional layer (Deterministic): Conv2d(6, 16, 5)
  • Linear layer with sparsity inducing prior + SELU activation function: Linear(16 $\times$ 5 $\times$ 5, 512)
  • Linear layer with sparsity inducing prior + SELU activation function: Linear(512, 256)
  • Linear layer with sparsity inducing prior + SELU activation function: Linear(256, 128)
  • Linear layer with sparsity inducing prior: Linear(128, 4)
Table 9 presents quantitative results for the variance of the weights and error metrics. The last column provides a measure of the sparsity-inducing effect of the Laplace prior on the linear layers.
Table 9: Bayesian neural network on CIFAR10 dataset. Test errors and log pointwise predictive density (LPPD) achieved using the final particle cloud with $N=50$. Computation times and standard deviation of the empirical distribution of the weight matrix $w$ for linear layers are also provided.
Algorithm Error (%) LPPD (×$\times 10^{-1}$) Time (s) Std. $w$
MYPGD $5.27\pm 0.95$ $-4.41\pm 0.38$ $201$ $3.10$
MYIPLA $\mathbf{5.23\pm 1.31}$ $-5.05\pm 0.45$ $199$ $3.22$
PIPGLA $5.39\pm 1.02$ $\mathbf{-4.32\pm 0.37}$ $295$ $\mathbf{2.85}$
PGD $6.01\pm 1.15$ $-5.73\pm 0.40$ $\mathbf{178}$ $11.51$
SOUL $9.11\pm 2.03$ $-7.68\pm 1.56$ $433$ $15.68$
IPLA $5.40\pm 1.33$ $-5.90\pm 0.75$ $181$ $15.73$
E.3.3 Non-Differentiable Activation Functions
During gradient checking in neural network training, a potential source of inaccuracy arises from the presence of non-differentiable points in the objective function [Kumar, 2024]. These non-smooth points often result from the use of activation functions such as the Rectified Linear Unit (ReLU), defined as $\max(0,x)$, as well as from the hinge loss in support vector machines, maxout neurons, among others. To give a concrete example, consider the ReLU activation function and $x<0$ but very close to 00. The analytic gradient evaluated at $x$ is equal to 00. However, the numerical gradient can be non-zero when using a finite difference approximation in case $x+h>0$.
Our method provides a principled way of dealing with these non-differentiable points. To illustrate this, we present a simple example similar to the one in the previous section. Here, we consider a Bayesian neural network with a Normal prior distribution on the weights to classify MNIST digits 1 and 7, instead of 4 and 9. Additionally, we use a linear approximation of $tanh$ as the activation function to mitigate the dying neuron problem associated with ReLU [Lu et al., 2020], while noting that it still remains non-differentiable. This linear approximation is defined as
$h(x)=\begin{cases}-1\quad\text{if}\;x<-1,\\ x\quad\ \ \;\text{if}\;x\in[-1,1],\\ 1\quad\ \;\;\text{if}\;x>1.\end{cases}$
Furthermore, we can compute the proximal mapping of $h$ which is given by
$\operatorname{prox}_{h}^{\lambda}(x)=\begin{cases}x&\text{if}\;x<-1,\\ -1&\text{if}\;x\in[-1,-1+\lambda],\\ x-\lambda&\text{for}\;x\in[-1+\lambda,1-\lambda],\\ 1&\text{for}\;x\in[1-\lambda,1],\\ x&\text{if}\;x>1,\end{cases}$ (53)
where we have applied the first order optimality condition and used the subgradient of the function at $x=-1$ and $1$ which is given by the sets $[0,1]$ and $[-1,0]$, respectively. Therefore, in this setting we have the following likelihood
$p(l|f,x)\propto\exp\bigg{(}\sum_{j=1}^{40}v_{lj}h_{LR}\Big{(}\sum_{i=1}^{784}w_{ji}f_{i}\Big{)}\bigg{)}.$
We assign priors $p_{\alpha}(w)=\prod_{i}\mathcal{N}(w_{i}|0,e^{2\alpha})$ and $p_{\beta}(v)=\prod_{i}\mathcal{N}(v_{i}|0,e^{2\beta})$ to the input and output layer’s weights, respectively, and learn $\theta=(\alpha,\beta)$ from the data. Hence, model’s density is given by
$p_{\theta}(x,\mathcal{Y}_{\text{train}})=\prod_{i}\mathcal{N}(w_{i}|0,e^{2\alpha})\prod_{j}\mathcal{N}(v_{j}|0,e^{2\beta})\prod_{(f,l)\in\mathcal{Y}_{\text{train}}}p(l|f,x),$
where $x$ denotes the weight matrices, i.e. $x=(w,v)$. We note that the log density can be decomposed as
$-\log p_{\theta}(x,\mathcal{Y}_{\text{train}})=\underbrace{d_{w}\alpha+\frac{1}{2}\sum_{i}|w_{i}|^{2}e^{-2\alpha}+d_{v}\beta+\frac{1}{2}\sum_{j}|v_{j}|^{2}e^{-2\beta}}_{g_{1}(\theta,x)}\underbrace{-\sum_{(f,l)\in\mathcal{Y}_{\text{train}}}\log p(l|f,x)}_{g_{2}(\theta,x)},$
where $d_{w}$ and $d_{v}$ denote the dimensions of the weights $w$ and $v$, respectively, $g_{1}$ is differentiable and depends on $\theta$ and $x$, while $g_{2}$ is proper, convex and lower semi-continuous and only depends on $x$, that is, $g_{2}(\theta,x)=g_{2}(x)$. As a result, the non-differentiability affects only the latent variables $x$. In this case, we can compute the proximity map of $g_{2}$ by using the expression for the proximity map of the activation function, provided in (53).
We follow the same implementation details as outlined in the previous section and use the same performance metrics: average classification error over a test set and log pointwise predictive density. The results for the proposed proximal algorithms are provided in Table 10 together with the computation times for $N=50$ and $500$ iterations. In addition, plots of the evolution of the different performance metrics for different number of particles are shown in Figure 8. We observe that the standard deviation of the LPPD across runs decreases as the number of particles increases. Moreover, PIPGLA exhibits a lower standard deviation compared to the other methods.
Table 10: Bayesian neural network with non-differentiable activation function. Test errors and log pointwise predictive density (LPPD) achieved using the final particle cloud with $N=50$ and $500$ iterations.
[…]
MYPGD $0.75\pm 0.68$ $\mathbf{-3.36\pm 1.18}$ $\mathbf{40}$
MYIPLA $\mathbf{0.70\pm 0.50}$ $-4.28\pm 2.86$ $\mathbf{40}$
PIPGLA $0.90\pm 0.49$ $-3.76\pm 0.96$ $68$
(a) Test error. $N=5$
(b) Test error. $N=50$
(c) Test error. $N=100$
(d) LPPD. $N=5$
(e) LPPD. $N=50$
(f) LPPD. $N=100$
Figure 8: Evolution of the classification error on a test set (top) and the log pointwise predictive density (LPPD) (bottom) over iterations in the BNN experiment with non-differentiable activation function, using $250$ iterations. Values averaged over 100 runs.
E.4 Image Deblurring
We consider the problem of recovering a high-quality image from a blurred and noisy observation $y=Hx+\varepsilon$, where $H$ is a blurring operator that blurs a pixel $x_{i,j}$ uniformly with its closest neighbours (10 $\times$ 10 patch), and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The log prior is proportional to the total variation defined as $TV(x)=\|\nabla_{d}x\|_{1}$, where $\|\cdot\|_{1}$ is the $\ell_{1}$ norm and $\nabla_{d}$ is the two-dimensional discrete gradient operator, which is non-differentiable. The proportionality parameter, $e^{\theta}$, which controls the strength of this log prior, typically requires manual tuning. Instead of fixing this parameter manually, we estimate its optimal value using our proposed algorithms. Note that we exponentiate $\theta$ to ensure its positivity.
The posterior distribution for the model takes the form
$p_{\theta}(y|x)\propto\exp\left(-\|y-Hx\|^{2}/(2\sigma^{2})-e^{\theta}TV(x)+\log C(\theta)\right),$
where $C(\theta)$ is proportional to the normalising constant of the prior distribution. To compute $C(\theta)$, we start by considering the case when $\theta=0$. In this case, the total variation prior is given by
$p(x)=C\exp(-TV(x)),$
where $C$ is constant. For $\theta\neq 0$, the prior $p_{\theta}(x)$ can be expressed using the pushforward measure as
$p_{\theta}(x)=T_{e^{\theta}}\#p(x)=e^{d_{x}\theta}p(e^{\theta}x),$
where $T_{e^{\theta}}\#$ denotes the pushforward operator and $d_{x}$ is the dimension of $x$. Due to the linearity of the total variation norm, it follows that
$p_{\theta}(x)=Ce^{d_{x}\theta}\exp\left(-e^{\theta}TV(x)\right).$
Thus, we obtain that theta $C(\theta)=e^{d_{x}\theta}$. For the experiments, we employ the algorithms proposed by [Douglas and Rachford, 1956] and [Chambolle, 2004] to efficiently compute the proximal operator of the total variation norm. Due to the difficulty of computing the joint proximal operator over the parameter $\theta$ and latent variables $x$, we have consider hybrid versions of the algorithms, which use standard gradient-based updates for the parameters and proximal updates for the particles. That is, the updates for the hybrid MYIPLA algorithm are given by
$\theta_{n+1}^{N}=$ $\theta_{n}^{N}-\frac{\gamma}{d_{x}N}\sum_{i=1}^{N}e^{\theta_{n}^{N}}TV(X_{n}^{i,N})+\gamma+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$
$X_{n+1}^{i,N}=$ $\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}^{i,N}-\gamma\frac{H^{\intercal}(HX_{n}^{i,N}-y)}{\sigma^{2}}+\frac{\gamma}{\lambda}\operatorname{prox}_{e^{\theta_{n}^{N}}TV}^{\lambda}(X_{n}^{i,N})+\sqrt{2\gamma}\;\xi_{n+1}^{i,N}.$
Note that as in the Bayesian neural network example, we apply the heuristic of dividing the gradient term in the $\theta$ updates by $d_{x}$ for numerical stability. For PIPGLA, the update for the parameter $\theta$ remains the same, while the updates for the particles are of the form
$X_{n+1/2}^{i,N}$ $=X_{n}^{i,N}-\gamma\frac{H^{\intercal}(HX_{n}^{i,N}-y)}{\sigma^{2}}+\sqrt{2\gamma}\;\xi_{n+1}^{i,N},$
$X_{n+1}^{i,N}$ $=\operatorname{prox}_{e^{\theta_{n+1}^{N}}TV}^{\lambda}\big{(}X_{n+1/2}^{i,N}\big{)}.$
Analogous forms are defined for the proximal PGD algorithms.
@@ E.4 Image Deblurring
We use black and white images with pixels values ranging from 0 to 255. The dimensions of the acoustic guitar image are $d_{x}=n_{1}\times n_{2}=584\times 238$, while the dimensions of the boat image (a standard benchmark in the image reconstruction literature) are $d_{x}=512\times 512$.
Implementation details.
We implement the proximal operator of the total variation using the proxTV Python package [Barbero and Sra, 2011, 2018]. Specifically, we employ the Douglas-Rachford method introduced by Douglas and Rachford [1956] and the Chambolle-Pock method [Chambolle, 2004]. The Douglas-Rachford method is significantly faster than the Chambolle-Pock method. It is important to note that increasing the precision of the Moreau-Yosida envelope significantly slows down the computation of the proximal operator when using these numerical schemes. We set $\gamma=0.01$, and $\lambda=0.4$ for MYPGD and MYIPLA and $\lambda=0.001$ for PIPGLA (after performing a grid search) which ensures that the algorithms are not close to losing stability. In addition, the pixels of the initial particles are drawn from a normal distribution with mean $\mu=50$ and scale parameter $10$, while the initial parameter estimate $\theta_{0}$ is sampled from a uniform distribution over $[-15,10]$.
Performance metrics.
To evaluate the performance of our algorithms in image reconstruction, we evaluate the mean squared error (MSE) and the structural similarity index (SSIM) between the particle cloud and the ground-truth image. The SSIM quantifies image quality by comparing luminance, contrast, and structural details.
(a) Original
(b) Blurred
(c) MYPGD
(d) MYIPLA
(e) PIPGLA
Figure 9: Image deblurring experiment. All the algorithms use $N=10$ particles and are run for 3000 iterations with a burn-in of 100 iterations.
(a) $\theta$ estimates
(b) MSE
(c) Structural similarity index (SSIM)
Figure 10: Evolution of different quantities over iterations in the image deblurring experiment with $N=10$ particles for the acoustic guitar image. The plots are shown after discarding a burn-in period of 100 iterations and the initial parameter is $\theta_{0}=-10.5$.
(a) Original
(b) Blurred
(c) MYPGD
(d) MYIPLA
(e) PIPGLA
Figure 11: Image deblurring experiment. All the algorithms use $N=10$ particles and are run for 3000 iterations with a burn-in of 100 iterations.
(a) $\theta$ estimates
(b) MSE
(c) Structural similarity index (SSIM)
Figure 12: Evolution of different quantities over iterations in the image deblurring experiment with $N=10$ particles for the boat image. The plots are shown after discarding a burn-in period of 100 iterations and the initial parameter is $\theta_{0}=-8.1$.
Results.
Figures 9 and 11 display the original and blurred images alongside the reconstructed images obtained using our different proximal algorithms. The methods are run for 3000 iterations (with a burn-in of 100 iterations) and $N=10$ particles, employing the Douglas-Rachford method to numerically evaluate the proximal operator of the total variation norm. Figures 10 and 12 illustrate the evolution of the parameter estimates $\theta$, the mean squared error and the SSIM, after discarding a burn-in period of 100 iterations. The high MSE for PIPGLA (Figures 10(b) and 12(b)) arises from the difference in the shades of grey between the reconstructed and the original images, remaining large regardless of the choice of the proximal parameter $\lambda$. Besides, the optimal value for the strength of the total variation prior achieved by the algorithms, $e^{\theta}\approx 0.35$ (for both test images) is close to the value set manually in similar works for image reconstruction (e.g. Pereyra [2016], Durmus et al. [2018], Goldman et al. [2022]).
E.5 Nuclear-norm models for low rank matrix estimation
In this section, we demonstrate another application of our methods: the problem of matrix completion. Matrix completion [Candès and Plan, 2010, Liu et al., 2018] focuses on recovering an intact matrix with low-rank property from incomplete data. Its application varies from wireless communications [Kortas et al., 2017], traffic sensing [Mardani and Giannakis, 2014] to integrated radar and recommender systems [Gogna and Majumdar, 2015]. The low-rank prior knowledge is incorporated in the model using the nuclear-norm of the matrix [Fazel, 2002]. However, similar to the image deblurring example, the strength of this prior is a hyperparameter that must be set manually. Instead, we estimate the optimal value of this parameter, thereby extending the applicability of proximal methods that typically perform MLE, rather than MMLE, as in our algorithms.
We conduct a graphical posterior predictive check of the widely used nuclear norm model for low-rank matrices, similar to the example in Pereyra [2016], but in the context of matrix completion rather than matrix denoising. Let $x$ be an unknown low-rank matrix of size $n_{1}\times n_{2}$. Consider a mask $M_{\Omega}$, where $\Omega$ is a set of indices from a matrix of size $n_{1}\times n_{2}$. When the mask is applied to the matrix $x$, i.e., $M_{\Omega}X$, only the entries of the matrix corresponding to indices in $\Omega$ are observed. Furthermore, after the masking operation, we do not have direct access to the observed entries but instead observe a noisy version of them, where the observational noise has mean zero and covariance $\sigma^{2}I$. Thus, our observations are given by $y=M_{\Omega}x+\sigma^{2}\varepsilon$, with $\varepsilon\sim\mathcal{N}(0,I)$. It is important to highlight, that we will also estimate with our algorithms the scale parameter $\sigma$, rather than requiring it to be fixed manually.
Our objective is to recover $x$ from $y$ under the prior knowledge that $x$ has low rank, that is, most of its singular values are zero. A convenient model for this type of problem is the nuclear norm prior, which is a sparsity-inducing prior, given by
$p_{\theta}(x)=C(\theta_{1})e^{-e^{\theta_{1}}\|x\|_{\text{tr}}},$
where $\|\cdot\|_{\text{tr}}$ is the trace (or nuclear) norm, which is a convex envelope of the rank function [Bach, 2008], and is defined as
$\|x\|_{\text{tr}}=\sum_{i=1}^{r}\sigma_{i}(x),$
$r=\text{rank}(x)$ and $\sigma_{1}(x)\geq\dots\geq\sigma_{r}(x)\geq 0$ are the singular values. Besides, the constant $C(\theta_{1})$ can be computed using the pushforward argument, as in Section E.4, leveraging the linearity of the trace norm. Specifically, it is given by $C(\theta_{1})=Ce^{d_{x}\theta_{1}}$, where $d_{x}$ denotes the dimension of the original matrix $x$ and $C$ is a constant. The posterior distribution of our model can be written as
$p_{\theta}(y|x)\propto\frac{e^{d_{x}\theta_{1}}}{e^{d_{y}\theta_{2}}}\exp\left({-\frac{\|M_{\Omega}x-y\|^{2}}{2e^{2\theta_{2}}}-e^{\theta_{1}}\|x\|_{\text{tr}}}\right).$
Therefore, the negative log density can be decomposed as
$U(\theta,x)=\underbrace{-d_{x}\theta_{1}+d_{y}\theta_{2}+\frac{\|M_{\Omega}x-y\|^{2}}{2e^{2\theta_{2}}}}_{g_{1}(\theta,x)}+\underbrace{e^{\theta_{1}}\|x\|_{\text{tr}}}_{g_{2}(\theta,X)},$
where $d_{y}$ denotes the number of observed entries. Note that we exponentiate the parameters $\theta_{1}$ and $\theta_{2}$ to ensure their positivity.
Dataset.
We use the checkerboard image of size $188\times 188$ and rank 2. We add Gaussian observational noise with variance $\sigma^{2}=0.1$ and mask $30\%$ of the pixels in the image.
@@ E.5 Nuclear-norm models for low rank matrix estimation
Recall that $g_{2}(\theta,x)$ is of the form
$g_{2}(\theta,x)=e^{\theta_{1}}\|x\|_{\text{tr}}.$
To compute the proximal map, we first observe that if $\theta_{1}$ is known, then by Cai et al. [2010, Theorem 2.1], it follows that
$\operatorname{prox}_{g_{2}}^{\lambda}(x)=\operatorname*{arg\,min}_{z}\;\{e^{\theta_{1}}\|z\|_{\text{tr}}+\frac{1}{2\lambda}\|x-z\|_{F}^{2}\}=S_{e^{\theta_{1}}\lambda}(x):=U\Sigma_{e^{\theta_{1}}\lambda}V^{T},$
where $U\Sigma V^{T}$ is a singular value decomposition, and $\Sigma_{\beta}$ is diagonal with entries $(\Sigma_{\beta})_{ii}=\max\{\Sigma_{ii}-\beta,0\}$. Based on this, we calculate
$\operatorname{prox}_{g_{2}}^{\lambda}(\theta,x)=\operatorname*{arg\,min}_{(\alpha,z)}\;\{e^{\alpha}\|z\|_{\text{tr}}+\frac{1}{2\lambda}\big{(}\|\theta_{1}-\alpha\|^{2}+\|x-z\|_{F}^{2}\big{)}\},$
[…]
$\alpha=\theta_{1}+\lambda e^{\alpha}\|S_{e^{\alpha}\lambda}(x)\|_{\text{tr}}\Longrightarrow(\alpha-\theta)e^{\theta_{1}-\alpha}=\lambda e^{\theta_{1}}\|S_{e^{\alpha}\lambda}(x)\|_{\text{tr}},$ (54)
$z=S_{\lambda e^{\alpha}}(x).$ (55)
@@
$(\alpha-\theta_{1})e^{\theta_{1}-\alpha}\approx\lambda e^{\theta_{1}}\|S_{e^{\theta}\lambda}(x)\|_{\text{tr}}\Longrightarrow\alpha\approx\theta_{1}+W(\lambda e^{\theta_{1}}\|S_{e^{\theta_{1}}\lambda}(X^{l})\|_{\text{tr}}),$
where $W$ is the Lambert $W$ function. Substituting this value of $\alpha$ into (55), we obtain
$z\approx S_{e^{\alpha}\lambda}(x).$
@@
To stabilise the implementation of the algorithms, we divide the gradient and proximal mapping terms in the updates of $\theta_{1}$ and $\theta_{2}$ by the dimension of the the matrix $x$, $d_{x}$, and the number of observed entries in $y$, $d_{y}$, respectively. We then set $\gamma=0.01$, and $\lambda=0.25$ for MYPGD and MYIPLA and $\lambda=0.01$ for PIPGLA. The pixels of the initial particles are drawn from a normal distribution with mean $\mu=50$ and scale parameter $10$, while the initial values of the parameters $\theta_{1}$ and $\theta_{2}$ are drawn from uniform distributions over $[-15,5]$ and $[-10,10]$, respectively.
Performance metrics.
To asses the performance of our algorithms for low-rank matrix completion, we analyse the normalised mean squared error (NMSE) for both the entire matrix and the missing entries.